Compensated Absences Matt Broderick Company began operations on January 2, 2016. The following information was provided concerning the companys employees who are paid hourly Number of individuals employed Number of hours worked per day Paid vacation days earned annually by each employee Paid sick days earned annually by each employee 9 employees 8 hours 10 days 6 days Vacation days may be taken after January 15 of the year following the year in which they are earned. Sick days may be taken as soon as they are earned, unused sick days accumulate. Additional information is as follows Actual Hourly Wage Rate Vacation Days Used by Each Emplo Sick Days Used by Each Emplo 2016 $10 2016 2017 $11 2016 4 2017 2017 Matt Broderick Company has chosen not to accrue paid sick leave until used, and has chosen to accrue vacation time at expected future rates of pay without discounting. The company used the follow ng projected rates to accrue vacation time Year in Which Vacation Time Was Earned 2016 2017 Projected Future Pay Rates Used to Accrue Vacation Pay $10.75 11.60asfroC (a) Preparejournal entries to record transactions related to compensated absences during 2016 and 2017 Accruals Rate 2016 10.75 10 7,740 2017 2017 Payment of Vacation Time Rate 2017 2017 2016 2016 2016 2017 2017 2017(b) Compute the amounts of any liability for compensated absences that should be reported on the balance sheet at December 31, 2016, and 2017 Accrued liability at year-end: 2016 2017 January 1 balance December 31 balance
Answer #1

Ans) a)

Journal Entries
for the year 2016
Year Transactions Debit Credit
2016 Salary & Wages Exp a/c 7,200 (i)
     To Salary & Wages Payable a/c 7,200(i)
(being accrue expenses for vacations)
2016 Salary & Wages Exp a/c 4320(ii)
        To Salary & Wages Payable a/c 4320(ii)
(being accrue expenses for sick leave pay)
2016 Salary & Wages Payable a/c 2880(iii)
    To cash a/c 2880(iii)
(being sick leave paid)

Calculations;

Formula-

Amount = Employees * Per hour * working hours per day * no of days

i) Amount= 9* $10*8*10= 7200

ii) Amount= 9* $10*8*6= 4320

iii) Amount= 9* $10*8*4= 2880

Journal Entries
for the year 2017
Year Transactions Debit Credit
2017 Salary & Wages Exp a/c 7920(i)
   To Salary & Wages Payable a/c 7920(i)
(being accrue the expenses for vacations)
2017 Salary & Wages Exp a/c 4752(ii)
     To Salary & Wages Payable a/c 4752(ii)
(being accure expenses for sick leave pay)
2017 Salary & Wages Exp a/c 648
Salary & Wages Payable a/c 6480(iii)
     To Cash a/c 7128(v)
(being payment of vacation period)
2017 Salary & Wages Exp a/c 144
Salary & Wages Payable a/c 3816(iv)
     To Cash a/c 3960(vi)
(being payment for sick leave)

Calculations;

Formula as stated above;

Amount= Employees * Per hour * working hours per day * no of days

i) amount= 9*$11*8*10=7920

ii) amount= 9*$11*8*6= 4752

iii) amount= 9*$10*8*9= 6480

iv) amount= (9*$11*8*3)2376 + 1440 (9*$10*8*2)=3816

v) amount= 9*$11*8*9= 7128

vi) amount= 9*$11*8*5= 3960

Note; amount for sick days and vacation paid at current wage. The FIFO (first in first out) basis has been applied whenever the payment needed to be done at different rates.

Ans b) 2016

Vacation wages Sick pay
payable wages payable
1-Jan $0 $0
add; accrued 7200 4320
less; paid -2880
Balance 7200(i) 1440(ii)

2017

Vacation wages Sick pay
payable wages payable
1-Jan $7,200 $1,440
add; accrued 7920 4752
less; paid -6480 -3816
Balance 8640(iii) 2376(iv)

Calculations;

Formula- Amount= Employees * Per hour * working hours per day * no of days

i) amount= 9*$10*8*10=7200

ii) amount= 9*$10*8*2=1440

iii) amount= (9*$10*8*1)720+ (9*$11*8*10)7920=8640

iv) amount= 9*$11*8*3= 2376